OpenStudy (anonymous):

simplify sec^2(theta) times cos^2(theta) - tan^2(theta)

OpenStudy (anonymous):

sec^2(theta) = 1/(cos^2(theta)) So sec^2(theta)*cos^2(theta) - tan^2(theta) = (1/(cos^2(theta)))*cos^2(theta) - tan^2(theta) = 1 - tan^2(theta)